# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ 6.2. Combinatorial forcing
This localised variant of combinatorial forcing for topological Ramsey spaces was introduced in [DMN15], along with various basic properties of this combinatorial forcing. The details of the proofs of these properties with our definition of semiselectivity (Definition 6.3) remains mostly unchanged, but we shall provide an overview for the sake of completion.
@@
    If $a\notin\mathcal{AR}\mathord{\upharpoonright}A$, $A$ accepts $a$.
@@
The proofs are straightforward and mostly the same as the non-localised variant of combinatorial forcing – we refer readers to Lemma 4.31 of [T10]. ∎
@@
By A4 mod $\mathcal{H}$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}$ or $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}^{c}$. It remains to show that the latter case is impossible. Since $A^{\prime}$ decides every $b\in\mathcal{AR}\mathord{\upharpoonright}A$, so does $B$. Therefore, if $r_{\operatorname{lh}(a)+1}[a,B]$ then $B$ accepts every $b\in r_{\operatorname{lh}(a)+1}[a,B]$. By Lemma 6.6(7), this implies that $B$ accepts $a$, contradicting that $A$ rejects $a$. ∎
@@
The abstract combinatorial forcing allows us to prove the abstract local Ellentuck theorem – see Theorem 3.12 of [DMN15]. We do not require the full strength of this result in this paper. Instead, we give an outline of the proof if $\mathcal{H}$ is semiselective, then every metrically open set $\mathcal{X}\subseteq\mathcal{R}$ is $\mathcal{H}$-Ramsey.
@@ 6.3. Abstract Mathias forcing
    If $A\in\mathcal{R}$, then $A_{G_{A}}=A$.
@@
We only prove (1) - (2) follows from the uniqueness of $A_{G}$ in (1), and (3) follow easily from the definitions. Let $\mathcal{D}_{n}:=\{(a,A)\in\mathbb{M}_{\mathcal{H}}:\operatorname{lh}(a)\geq n\}$. We see that $\mathcal{D}_{n}$ is dense open in $\mathbb{M}_{\mathcal{H}}$ for all $n$ - if $(a,A)\notin\mathcal{D}_{n}$ (i.e. $\operatorname{lh}(a)<n$) and $b\in r_{n}[a,A]$, then by A3 mod $\mathcal{H}$ we have that $b,A]\cap\mathcal{H}\neq\emptyset$. Let $B\in[b,A]\cap\mathcal{H}$, and we have that $b,B]\subseteq[a,A]$, so $(b,B)\leq(a,A)$ and $(b,B)\in\mathcal{D}_{n}$. Therefore, $G\cap\mathcal{D}_{n}\neq\emptyset$ for all $n$, and since $G$ is a filter, for all $(a,A),(b,B)\in G$, $a\sqsubseteq b$ or $b\sqsubseteq a$. Thus, there exists a unique $A_{G}\in\mathcal{R}$ such that for all $n$, $(r_{n}(A_{G}),A)\in G$ for some $A$. It’s easy to verify that $G=G_{A_{G}}$. ∎
@@
Let $A_{G}$ be an $\mathbb{M}_{\mathcal{H}}$-generic element of $\mathcal{R}$ (over some ground model $\mathbf{V}$), and let $A_{H}\leq A_{G}$. We need to show that the filter:
$H:=\{(a,A)\in\mathbb{M}_{\mathcal{H}}:A_{H}\in[a,A]\}$
@@
It’s clear that $\mathcal{E}_{a}$ is open, so shall show that it is dense in $\mathcal{H}\mathord{\upharpoonright}[a,A]$. By the comment above, let $A^{\prime}\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a,A^{\prime}]\subseteq\overline{\mathcal{D}}$. Let $\{b_{i}:1\leq i\leq N\}$ be an enumeration of all $b\preceq_{\mathrm{fin}}a$. Let $A_{0}:=A^{\prime}$, and suppose that we have defined $A_{i}\in[a,A^{\prime}]$. We have that $b_{i+1}\in\mathcal{AR}\mathord{\upharpoonright}A_{i}$ (as $b_{i+1}\leq_{\mathrm{fin}}a\sqsubseteq A_{i}$), so by the observation stated prior to the claim there exists some $A_{i+1}^{\prime}\in\mathcal{H}\mathord{\upharpoonright}[b_{i+1},A_{i}]$ which accepts $b_{i+1},A_{i}$ w.r.t. $\overline{\mathcal{D}}$, i.e. $b_{i+1},A_{i+1}^{\prime}]\subseteq\overline{\mathcal{D}}$. By A3 mod $\mathcal{H}$, there exists some $A_{i+1}\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a_{i+1},A_{i+1}]\subseteq[a_{i+1},A_{i+1}^{\prime}]$. This completes the induction, and $(a,A_{N})\in\mathcal{D}^{\prime}$, as desired. ∎
Fix any $(a,A)\in G$. Since $\mathcal{H}$ is semiselective, for all $(b,B)\leq(a,A)$ in $\mathbb{M}_{\mathcal{H}}$, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}[b,B]$ which diagonalises $\vec{\mathcal{E}}_{b}:=\{\mathcal{E}_{c}\}_{c\in\mathcal{AR}\mathord{\upharpoonright}[b,A]}$ within $A$. That is, the set:
$\mathcal{D}^{\prime}:=\{(b,B)\in\mathbb{M}_{\mathcal{H}}:\text{$B$ diagonalises $\vec{\mathcal{E}}_{b}$ within $A$}\}$
is dense open in $\mathbb{M}_{\mathcal{H}}$. Since $G$ is generic, we may let $(b,B)\in G\cap\mathcal{D}^{\prime}$. Then $A_{G}\in[b,B]$, and for all $c\in\mathcal{AR}\mathord{\upharpoonright}[b,A_{G}]$ and $d\preceq_{\mathrm{fin}}c$, $d,A_{G}]\subseteq[d,B]\subseteq\overline{\mathcal{D}}$. Let $n$ be the least integer such that $0pt_{A_{G}}(r_{n}(A_{H}))=:m\geq\operatorname{lh}(b)$. Then $r_{n}(A_{H})\preceq_{\mathrm{fin}}r_{m}(A_{G})\in\mathcal{AR}\mathord{\upharpoonright}[b,B]$, so $A_{H}\in[r_{n}(A_{H}),A_{G}]\subseteq[r_{n}(A_{H}),B]\subseteq\overline{\mathcal{D}}$. This implies that $(r_{n}(A_{H}),B)\in H\cap\mathcal{D}$, and since $\mathcal{D}$ is arbitrary, $H$ is a generic filter as required. ∎
@@ 7. Topological Ramsey space of block subspaces
By Hindman’s theorem, the pigeonhole principle hold for $E^{[\infty]}$ if $|\mathbb{F}|=2$. That is, for all $Y\subseteq E$ and $A\in E^{[\infty]}$, there exists some $B\leq A$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$. Consequently, if for each $A:=(x_{n})_{n<\omega}\in E^{[\infty]}$ we let:
@@
then $(E^{[\infty]},\leq,r)$ is a closed triple satisfying A1-A4. We shall discuss the intersection of the local Ramsey theory of $(E^{[\infty]},\leq,r)$ as a topological Ramsey space, and the local Ramsey theory of block subspaces developed Smythe, followed by an application of the abstract Mathias forcing to prove Theorem 1.7.
@@ 7.1. Semicoideals of block subspaces
On one hand, if $A\in\mathcal{H}$ and $a,A]\subseteq[a,B]$ for some $a\in\mathcal{AR}\mathord{\upharpoonright}A$, then by A3 mod $\mathcal{H}$ there exists some $C\in[a,A]\cap\mathcal{H}$. Then $C\in[a,B]$, and in particular $C\leq B$, so $B\in\mathcal{H}$ by $\leq$-upward closure. Conversely, suppose that $\mathcal{H}$ is $\leq^{*}$-upward closed. Let $A\in\mathcal{H}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$.
  1. (1)
@@
    Is Definition 6.3 equivalent to the natural extension of semiselectivity of coideal of $\omega]^{\omega}$ to semicoideal of block subspaces (i.e. Lemma 7.10, [T10])?
@@
    Is a selective (semi)coideal semiselective?
@@
In this statement (2) of the lemma, we also say that $B$ diagonalises $\{\mathcal{D}_{n}\}_{n<\omega}$ within $A$.
@@
It’s clear that $\mathcal{D}_{b}\subseteq\mathcal{H}\mathord{\upharpoonright}[b,A]$ for all $b\in\mathcal{AR}\mathord{\upharpoonright}A$, and that $\mathcal{D}_{b}$ is $\leq$-downward closed in $\mathcal{H}\mathord{\upharpoonright}[b,A]$. We see that $\mathcal{D}_{b}$ is dense in $\mathcal{H}\mathord{\upharpoonright}[b,A]$ – if $B\in\mathcal{H}\mathord{\upharpoonright}[b,A]$, then since $\mathcal{D}_{d_{A}(b)}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}A$, there exists some $C\in\mathcal{D}_{d_{A}(b)}$ such that $C\leq B$. Then $b^{\frown}(C/b)\in\mathcal{D}_{b}$ and $b^{\frown}(C/b)\leq B$. Since $\mathcal{H}$ is semiselective, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that for all $b\in\mathcal{AR}\mathord{\upharpoonright}B$, $b,B]\subseteq[b,A_{b}]$ for some $A_{b}\in\mathcal{D}_{b}$. In particular, $B\in[r_{n}(B),B]\subseteq[r_{n}(B),A_{r_{n}(B)}]$ for some $A_{r_{n}(B)}\in\mathcal{D}_{r_{n}(B)}$, so $B/n\leq A_{r_{n}(B)}/r_{n}(B)\in\mathcal{D}_{d_{A}(r_{n}(B))}$.
@@
An abstract notion of selectivity was also proposed in [DMN15], but the definition appears to be problematic. For instance, Proposition 5.4 of [DMN15] claims that every selective coideal is semiselective, but there appears to be a gap in the proof – it is unclear how the element $A^{2,1}\in\mathcal{D}_{a_{1}^{2}}\mathord{\upharpoonright}A^{1,n_{1}}$ may be obtained. It seems unlikely that this version of selectivity coincides with our definition of selectivity.
@@
We shall dedicate the remaining portion of this subsection to showing that if $\mathcal{U}\subseteq^{[\infty]}$ is a selective ultrafilter, then a block sequence $A_{G}\in(E^{[\infty]})^{\mathbf{V}[G]}$ is $\mathbb{M}_{\mathcal{U}}$-generic iff $A_{G}\leq^{*}A$ for all $A\in\mathcal{U}$. The proof is very similar to the case for the Ellentuck space – a proof may be found in, for instance, [H22].
The following proposition, in the setting of topological Ramsey spaces, was stated as Theorem 6.6 of [DMN15], extending the original theorem in the setting $\omega]^{\omega}$ in [F98]. We shall reprove this result for block subspaces.
@@
Assume that $|\mathbb{F}|=2$. Let $\mathcal{H}\subseteq E^{[\infty]}$ be a coideal. The following are equivalent:
@@
$\mathcal{E}:=\{B\in\mathcal{H}\mathord{\upharpoonright}A:\text{$B$ diagonalises $\{\mathcal{D}_{n}\}_{n<\omega}$ within $A$}\}$
is dense open below $A$ in $\mathcal{H}$. By the genericity of $\mathcal{U}$, let $B\in\mathcal{U}\cap\mathcal{E}$. Then $B/n\in\mathcal{D}_{d_{A}(r_{n}(B))}$ for all $n$, and since $\mathcal{U}$ is a filter, $B$ is compatible with $A_{n}$ for all $n$ in $\mathcal{U}$, so $B/n\leq A_{d_{A}(r_{n}(B))}$ for all $n$. Thus, $B$ diagonalises $(A_{n})_{n<\omega}$ within $A$.
@@
We remark that if $\mathcal{AR}$ is countable (e.g. $E^{[<\infty]}$ is countable), then the statement “$(a,A)$ captures $\mathcal{D}$” is $\Pi_{1}$ with respect to the Polish topology of $\mathcal{R}$. Since $\mathcal{R}$ is a subspace of $\mathcal{AR}^{\mathbb{N}}$ and $\mathcal{AR}$ is equipped with the discrete topology, we may apply Mostowki’s absoluteness th55eorem to conclude that “$(a,A)$ captures $\mathcal{D}$” is absolute.
@@
Assume that $|\mathbb{F}|=2$. Let $\mathcal{U}\subseteq E^{[\infty]}$ be a selective ultrafilter. Let $\mathcal{D}\subseteq\mathbb{M}_{\mathcal{U}}$ be dense open. For all $a\in E^{[<\infty]}$ and $A\in E^{[\infty]}$ such that $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, there exists some $B\in\mathcal{U}\mathord{\upharpoonright}[a,A]$ such that $(a,A)$ captures $\mathcal{D}$.
@@
For each $b\in E^{[<\infty]}\mathord{\upharpoonright}[a,A]$, let $A_{b}\in\mathcal{U}\mathord{\upharpoonright}[b,A]$ be any block sequence such that $(b,A_{b})\in\mathcal{D}$ if it exists – otherwise, we let $A_{b}:=A$. Since $\mathcal{U}$ is an ultrafilter, we may inductively define a $\leq$-decreasing sequence in $\mathcal{U}\mathord{\upharpoonright}[a,A]$ such that $A_{n}\leq A_{b}/a$ for all $b\in E^{[<\infty]}\mathord{\upharpoonright}[a,A]$ in which $d_{A}(b)\leq n$. Since $\mathcal{U}$ is selective, there exists some $B^{\prime}\in\mathcal{H}\mathord{\upharpoonright}(A/a)$ which diagonalises $(A_{n})_{n<\omega}$ within $A$. Let $B:=a^{\frown}B^{\prime}\in\mathcal{U}\mathord{\upharpoonright}[a,A]$, and consider the set:
@@
Since $\mathcal{X}$ is a metrically open subset of $E^{[<\infty]}$, by Proposition 6.11 there exists some $C\in\mathcal{U}\mathord{\upharpoonright}[a,B]$ such that $a,C]\subseteq\mathcal{X}$ or $a,C]\subseteq\mathcal{X}^{c}$. Observe that if $a,C]\subseteq\mathcal{X}$, then $(a,C)$ captures $\mathcal{D}$, so we’re done. It remains to show that $a,C]\subseteq\mathcal{X}^{c}$ is not possible. Since $\mathcal{D}$ is dense in $\mathbb{M}_{\mathcal{H}}$, there exists some condition $(b,D)\leq(a,C)$ in $\mathcal{D}$. This means that $(b,A_{b})\in\mathcal{D}$, and since $B/b\leq A_{d_{A}(b)}\leq A_{b}$, we have that $(b,B)\in\mathcal{D}$. In other words, there exists some $b\in E^{[<\infty]}\mathord{\upharpoonright}C$ such that $(b,B)\in\mathcal{D}$, so $C\notin\mathcal{X}$. ∎
@@
Assume that $|\mathbb{F}|=2$. Let $\mathcal{U}\subseteq E^{[\infty]}$ be a selective ultrafilter. Then a block sequence $A_{G}$ is $\mathbb{M}_{\mathcal{U}}$-generic iff $A_{G}\leq^{*}A$ for all $A\in\mathcal{U}$.
@@
Given a block sequence $A_{G}$, we let:
$G:=\{(a,A)\in\mathbb{M}_{\mathcal{U}}:A_{G}\in[a,A]\}.$
$\implies$: Suppose that $A_{G}$ is $\mathbb{M}_{\mathcal{U}}$-generic, i.e. $G$ is a generic filter. For any $A\in\mathcal{U}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, the set:
@@
is dense open in $\mathbb{M}_{\mathcal{U}}$ – indeed, for any $(b,B)\in\mathbb{M}_{\mathcal{U}}$, let $C\in\mathcal{U}$ be such that $C\leq B$ and $C\leq A$. Then $(a,a^{\frown}(C/a))\in\mathcal{D}_{A}$, and $C\leq^{*}B$ and $C\leq^{*}B$. Since $G$ is generic, $G\cap\mathcal{D}_{A}\neq\emptyset$, so we may let $(b,B)\in\mathcal{D}_{A}$ such that $A_{G}\in[b,B]$. This implies that $A_{G}\leq^{*}B\leq^{*}A$, as desired.

### v5

@@ 6.2. Combinatorial forcing
This localised variant of combinatorial forcing for topological Ramsey spaces was introduced in [6], along with various basic properties of this combinatorial forcing. The details of the proofs of these properties with our definition of semiselectivity (Definition 6.3) remains mostly unchanged, but we shall provide an overview for the sake of completion.
@@
    If $a\notin\mathcal{AR}\mathord{\upharpoonright}A$, $A$ accepts $a$ (as $a,A]=\emptyset$).
@@
The proofs are straightforward and mostly the same as the non-localised variant of combinatorial forcing — we refer readers to Lemma 4.31 of [27]. ∎
@@
By A4 mod $\mathcal{H}$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}$ or $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}^{c}$. It remains to show that the latter case is impossible. Since $A^{\prime}$ decides every $b\in\mathcal{AR}\mathord{\upharpoonright}A$, so does $B$. Therefore, if $r_{\operatorname{lh}(a)+1}[a,B]\subseteq\mathcal{O}^{c}$ then $B$ accepts every $b\in r_{\operatorname{lh}(a)+1}[a,B]$. By Lemma 6.6(7), this implies that $B$ accepts $a$, contradicting that $A$ rejects $a$. ∎
@@
The abstract combinatorial forcing allows us to prove the abstract local Ellentuck theorem — see Theorem 3.12 of [6]. We do not require the full strength of this result in this paper. Instead, we give an outline of the proof if $\mathcal{H}$ is semiselective, then every metrically open set $\mathcal{X}\subseteq\mathcal{R}$ is $\mathcal{H}$-Ramsey.
@@ 6.3. Abstract Mathias forcing
    If $A\in\mathcal{R}$ is $M_{\mathcal{H}}$-generic, then $A_{G_{A}}=A$.
@@
We only prove (1) — (2) follows from the uniqueness of $A_{G}$ in (1), and (3) follow easily from the definitions. Let $\mathcal{D}_{n}:=\{(a,A)\in\mathbb{M}_{\mathcal{H}}:\operatorname{lh}(a)\geq n\}$. We see that $\mathcal{D}_{n}$ is dense open in $\mathbb{M}_{\mathcal{H}}$ for all $n$ — if $(a,A)\notin\mathcal{D}_{n}$ (i.e. $\operatorname{lh}(a)<n$) and $b\in r_{n}[a,A]$, then by A3 mod $\mathcal{H}$ we have that $b,A]\cap\mathcal{H}\neq\emptyset$. Let $B\in[b,A]\cap\mathcal{H}$, and we have that $b,B]\subseteq[a,A]$, so $(b,B)\leq(a,A)$ and $(b,B)\in\mathcal{D}_{n}$. Therefore, $G\cap\mathcal{D}_{n}\neq\emptyset$ for all $n$, and since $G$ is a filter, for all $(a,A),(b,B)\in G$, $a\sqsubseteq b$ or $b\sqsubseteq a$. Thus, there exists a unique $A_{G}\in\mathcal{R}$ such that for all $n$, $(r_{n}(A_{G}),A)\in G$ for some $A$. It’s easy to verify that $G=G_{A_{G}}$. ∎
@@
Let $A_{G}$ be an $\mathbb{M}_{\mathcal{H}}$-generic element of $\mathcal{R}$ (over some ground model $\mathbf{V}$), and let $D\leq A_{G}$. We need to show that the filter:
$H:=\{(a,A)\in\mathbb{M}_{\mathcal{H}}:D\in[a,A]\}$
@@
It’s clear that $\mathcal{E}_{a}$ is open, so we shall show that it is dense in $\mathcal{H}\mathord{\upharpoonright}[a,A]$. By the comment above, let $A^{\prime}\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a,A^{\prime}]\subseteq\overline{\mathcal{D}}$. Let $\{b_{i}:1\leq i\leq N\}$ be an enumeration of all $b\preceq_{\mathrm{fin}}a$. Let $A_{0}:=A^{\prime}$, and suppose that we have defined $A_{i}\in[a,A^{\prime}]$. We have that $b_{i+1}\in\mathcal{AR}\mathord{\upharpoonright}A_{i}$ (as $b_{i+1}\leq_{\mathrm{fin}}a\sqsubseteq A_{i}$), so by the observation stated prior to the claim there exists some $A_{i+1}^{\prime}\in\mathcal{H}\mathord{\upharpoonright}[b_{i+1},A_{i}]$ which accepts $b_{i+1},A_{i}$ w.r.t. $\overline{\mathcal{D}}$, i.e. $b_{i+1},A_{i+1}^{\prime}]\subseteq\overline{\mathcal{D}}$. By A3 mod $\mathcal{H}$, there exists some $A_{i+1}\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a_{i+1},A_{i+1}]\subseteq[a_{i+1},A_{i+1}^{\prime}]$. This completes the induction, and $(a,A_{N})\in\mathcal{D}$, as desired. ∎
Fix any $(a,A)\in G$. Since $\mathcal{H}$ is semiselective, for all $(b,B)\leq(a,A)$ in $\mathbb{M}_{\mathcal{H}}$, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}[b,B]$ which diagonalises $\vec{\mathcal{E}}_{b}:=\{\mathcal{E}_{c}\}_{c\in\mathcal{AR}\mathord{\upharpoonright}[b,A]}$ below $A$. That is, the set:
$\overline{\mathcal{D}}:=\{(b,B)\in\mathbb{M}_{\mathcal{H}}:\text{$B$ diagonalises $\vec{\mathcal{E}}_{b}$ below $A$}\}$
is dense open in $\mathbb{M}_{\mathcal{H}}$. Since $G$ is generic, we may let $(b,B)\in G\cap\mathcal{D}^{\prime}$. Then $A_{G}\in[b,B]$, and for all $c\in\mathcal{AR}\mathord{\upharpoonright}[b,A_{G}]$ and $d\preceq_{\mathrm{fin}}c$, $d,A_{G}]\subseteq[d,B]\subseteq\overline{\mathcal{D}}$. Let $n$ be the least integer such that $0pt_{A_{G}}(r_{n}(D))=:m\geq\operatorname{lh}(b)$. Then $r_{n}(D)\preceq_{\mathrm{fin}}r_{m}(A_{G})\in\mathcal{AR}\mathord{\upharpoonright}[b,B]$, so $D\in[r_{n}(D),A_{G}]\subseteq[r_{n}(D),B]\subseteq\overline{\mathcal{D}}$. This implies that $(r_{n}(D),B)\in H\cap\mathcal{D}$, and since $\mathcal{D}$ is arbitrary, $H$ is a generic filter as required. ∎
@@ 7. Topological Ramsey space of block subspaces
By Hindman’s theorem, the pigeonhole principle holds for $E^{[\infty]}$ if $E$ is a vector space over $\mathbb{F}_{2}$. That is, for all $Y\subseteq E$ and $A\in E^{[\infty]}$, there exists some $B\leq A$ such that $\langle B\rangle\subseteq Y$ or $\langle B\rangle\subseteq Y^{c}$. Consequently, if for each $A:=(x_{n})_{n<\omega}\in E^{[\infty]}$ we let:
@@
then $(E^{[\infty]},\leq,r)$ is a closed triple satisfying A1-A4. We shall discuss the intersection between the local Ramsey theory of $(E^{[\infty]},\leq,r)$ as a topological Ramsey space, and the local Ramsey theory of block subspaces developed Smythe, followed by an application of the abstract Mathias forcing to prove Theorem 1.7.
The symbols << and $\leq$ will appear repeatedly throughout this section. To avoid confusion, the symbol << shall be reserved for Notation 2.2(5) and (6), while $\leq$ shall be reserved for the quasi-order $\leq$ in $(E^{[\infty]},\leq,r)$.
@@ 7.1. Semicoideals of block subspaces
[…]
  1. (a)
@@
    Is Definition 6.3 equivalent to the natural extension of semiselectivity of coideal of $\omega]^{\omega}$ to semicoideal of block subspaces (i.e. Lemma 7.10, [27])?
@@
    Is a selective (semi)coideal (as in Definition 3.8) semiselective?
@@
In this statement (2) of the lemma, we also say that $B$ diagonalises $\{\mathcal{D}_{n}\}_{n<\omega}$ below $A$.
@@
It’s clear that $\mathcal{D}_{b}\subseteq\mathcal{H}\mathord{\upharpoonright}[b,A]$ for all $b\in\mathcal{AR}\mathord{\upharpoonright}A$, and that $\mathcal{D}_{b}$ is $\leq$-downward closed in $\mathcal{H}\mathord{\upharpoonright}[b,A]$. We see that $\mathcal{D}_{b}$ is dense in $\mathcal{H}\mathord{\upharpoonright}[b,A]$ — if $B\in\mathcal{H}\mathord{\upharpoonright}[b,A]$, then since $\mathcal{D}_{d_{A}(b)}$ is dense open in $\mathcal{H}\mathord{\upharpoonright}A$, there exists some $C\in\mathcal{D}_{d_{A}(b)}$ such that $C\leq B$. Then $b^{\frown}(C/b)\in\mathcal{D}_{b}$ and $b^{\frown}(C/b)\leq B$. Since $\mathcal{H}$ is semiselective, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ such that for all $b\in\mathcal{AR}\mathord{\upharpoonright}B$, $b,B]\subseteq[b,A_{b}]$ for some $A_{b}\in\mathcal{D}_{b}$. In particular, for all $n$ we have that $B\in[r_{n}(B),B]\subseteq[r_{n}(B),A_{r_{n}(B)}]$ for some $A_{r_{n}(B)}\in\mathcal{D}_{r_{n}(B)}$, so $B/n\leq A_{r_{n}(B)}/r_{n}(B)\in\mathcal{D}_{d_{A}(r_{n}(B))}$.
@@
An abstract notion of selectivity was also proposed in [6], but the definition appears to be problematic. For instance, Proposition 5.4 of [6] claims that every selective coideal is semiselective, but there appears to be a gap in the proof — it is unclear how the element $A^{2,1}\in\mathcal{D}_{a_{1}^{2}}\mathord{\upharpoonright}A^{1,n_{1}}$ may be obtained. It seems unlikely that this version of selectivity coincides with our definition of selectivity.
@@
We shall dedicate the remaining portion of this subsection to showing that if $\mathcal{U}\subseteq E^{[\infty]}$ is a selective ultrafilter, then a block sequence $D\in(E^{[\infty]})^{\mathbf{V}[G]}$ is $\mathbb{M}_{\mathcal{U}}$-generic iff $D\leq^{*}A$ for all $A\in\mathcal{U}$. The proof is very similar to the case for the Ellentuck space — a proof may be found in, for instance, Chapter 26 of [9].
The following proposition, in the setting of topological Ramsey spaces, was stated as Theorem 6.6 of [6], extending the original theorem in the setting $\omega]^{\omega}$ in [7]. We shall reprove this result for block subspaces.
@@
Suppose that $E$ is a vector space over $\mathbb{F}_{2}$. Let $\mathcal{H}\subseteq E^{[\infty]}$ be a coideal. The following are equivalent:
@@
$\mathcal{E}:=\{B\in\mathcal{H}\mathord{\upharpoonright}A:\text{$B$ diagonalises $\{\mathcal{D}_{n}\}_{n<\omega}$ below $A$}\}$
is dense open below $A$ in $\mathcal{H}$. By the genericity of $\mathcal{U}$, let $B\in\mathcal{U}\cap\mathcal{E}$. Then $B/n\in\mathcal{D}_{d_{A}(r_{n}(B))}$ for all $n$, and since $\mathcal{U}$ is a filter, $B$ is compatible with $A_{n}$ for all $n$ in $\mathcal{U}$, so $B/n\leq A_{d_{A}(r_{n}(B))}$ for all $n$. Thus, $B$ diagonalises $(A_{n})_{n<\omega}$ below $A$.
@@
We remark that if $\mathcal{AR}$ is countable (e.g. $E^{[<\infty]}$ is countable), then the statement “$(a,A)$ captures $\mathcal{D}$” is $\Pi_{1}$ with respect to the Polish topology of $\mathcal{R}$. Since $\mathcal{R}$ is a subspace of $\mathcal{AR}^{\mathbb{N}}$ and $\mathcal{AR}$ is equipped with the discrete topology, we may apply Mostowki’s absoluteness theorem to conclude that “$(a,A)$ captures $\mathcal{D}$” is absolute.
@@
Suppose that $E$ is a vector space over $\mathbb{F}_{2}$. Let $\mathcal{U}\subseteq E^{[\infty]}$ be a selective ultrafilter. Let $\mathcal{D}\subseteq\mathbb{M}_{\mathcal{U}}$ be dense open. For all $a\in E^{[<\infty]}$ and $A\in E^{[\infty]}$ such that $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, there exists some $B\in\mathcal{U}\mathord{\upharpoonright}[a,A]$ such that $(a,B)$ captures $\mathcal{D}$.
@@
For each $b\in E^{[<\infty]}\mathord{\upharpoonright}[a,A]$, let $A_{b}\in\mathcal{U}\mathord{\upharpoonright}[b,A]$ be any block sequence such that $(b,A_{b})\in\mathcal{D}$ if it exists — otherwise, we let $A_{b}:=A$. Since $\mathcal{U}$ is an ultrafilter, we may inductively define a $\leq$-decreasing sequence in $\mathcal{U}\mathord{\upharpoonright}[a,A]$ such that $A_{n}\leq A_{b}/a$ for all $b\in E^{[<\infty]}\mathord{\upharpoonright}[a,A]$ in which $d_{A}(b)\leq n$. Since $\mathcal{U}$ is selective, there exists some $B^{\prime}\in\mathcal{H}\mathord{\upharpoonright}(A/a)$ which diagonalises $(A_{n})_{n<\omega}$ below $A$. Let $B:=a^{\frown}B^{\prime}\in\mathcal{U}\mathord{\upharpoonright}[a,A]$, and consider the set:
@@
Since $\mathcal{X}$ is a metrically open subset of $E^{[\infty]}$, by Proposition 6.11 there exists some $C\in\mathcal{U}\mathord{\upharpoonright}[a,B]$ such that $a,C]\subseteq\mathcal{X}$ or $a,C]\subseteq\mathcal{X}^{c}$. Observe that if $a,C]\subseteq\mathcal{X}$, then $(a,C)$ captures $\mathcal{D}$, so we’re done. It remains to show that $a,C]\subseteq\mathcal{X}^{c}$ is not possible. Since $\mathcal{D}$ is dense in $\mathbb{M}_{\mathcal{H}}$, there exists some condition $(b,D)\leq(a,C)$ in $\mathcal{D}$. This means that $(b,A_{b})\in\mathcal{D}$, and since $B/b\leq A_{d_{A}(b)}\leq A_{b}$, we have that $(b,B)\in\mathcal{D}$. In other words, there exists some $b\in E^{[<\infty]}\mathord{\upharpoonright}C$ such that $(b,B)\in\mathcal{D}$, so $C\in\mathcal{X}$. ∎
@@
Suppose that $E$ is a vector space over $\mathbb{F}_{2}$. Let $\mathcal{U}\subseteq E^{[\infty]}$ be a selective ultrafilter. Then a block sequence $D$ is $\mathbb{M}_{\mathcal{U}}$-generic iff $D\leq^{*}A$ for all $A\in\mathcal{U}$.
@@
Given a block sequence $D$, we let:
$G:=\{(a,A)\in\mathbb{M}_{\mathcal{U}}:D\in[a,A]\}.$
$\implies$: Suppose that $D$ is $\mathbb{M}_{\mathcal{U}}$-generic, i.e. $G$ is a generic filter. For any $A\in\mathcal{U}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, the set:
@@
is dense open in $\mathbb{M}_{\mathcal{U}}$ — indeed, for any $(b,B)\in\mathbb{M}_{\mathcal{U}}$, let $C\in\mathcal{U}$ be such that $C\leq B$ and $C\leq A$. Then $(a,a^{\frown}(C/a))\in\mathcal{D}_{A}$, and $C\leq^{*}B$ and $C\leq^{*}B$. Since $G$ is generic, $G\cap\mathcal{D}_{A}\neq\emptyset$, so we may let $(b,B)\in\mathcal{D}_{A}$ such that $D\in[b,B]$. This implies that $D\leq^{*}B\leq^{*}A$, as desired.